To debug the issue of the suspected drifting TTs further, I temporarily hijacked CH0-CH8 of ADC1 in the c1lsc expansion chassis, and connected the "MON" outputs of the coil drivers (D010001) to them via some DB9 breakouts. The idea is to see if the problem is electrical. We should see some slow drift in the voltage to the TTs correlated with the spot walking off the IPPOS QPD. From the wiring diagram, it doesn't look like there is any monitoring (slow or fast) of the control voltages to the TT coils, this should be factored into the Acromag upgrade of c1iscaux/c1iscaux2. EPICS monitoring should be sufficient for this purpose so I didn't setup any new DQ channels, I'll just look at the EPICS from the IOP model.